Solve the equation 1 - 2/x + 1 = 2x/x + 1. Check the solution.
Answer by mananth(16949)

1 - 2/x + 1 = 2x /x + 1
-1 / x+1 = 2x/ x+1
-1(x+1)=2x(x+1)
-x-1=2x^2+2x
2x^2+3x+1=0
2x^2+2x+x+1=0
2x(x+1)+1(x+1)=0
(2x+1)(x+1)=0
x=-1/2 Or x=-1
..
CHECK
-1 / (-1/2+1) = 2*(-1/2)/ -1/2 +1
-2 = -2